# Improving Offline-to-Online Reinforcement Learning with Q-Ensembles

### v3

@@ 4.2 Loosing Pessimism
Specifically, we compare several Q-target computation methods. (a) MinQ is what we use in OnlineRL-N, where the minimum value of all the Q-networks in the ensemble is taken to compute the target. (b) MeanQ leverages the average of all the Q-values to compute the target. (c) REM is a method originally proposed to boost performance of DQN in the discrete-action setting, which uses the random convex combination of Q-values to compute the target (Agarwal et al., 2020). It is similar to ensemble average (MeanQ), but with more randomization. (d) RandomMinPair uses a minimization over a random subset 2 of the $N$ Q-functions, which is proposed in prior methods (Chen et al., 2021). (e) WeightedMinPair computes the target as the expectation of all the RandomMinPair targets, where the expectation is taken over all $N$-choose-2 pairs of Q-functions. RandomMinPair can be considered as a uniform-sampled version of WeightedMinPair.
Figure 3 presents the results of using different Q-target computation methods in the online phase based on OnlineRL-N. With MinQ, which is originally used in OnlineRL-N, as the bound, both MeanQ and REM exhibit poor performance, while RandomMinPair and WeightedMinPair outperform the other candidates with their efficient and stable online learning process. As the WeightedMinPair method is more stable on many datasets than the RandomMinPair method, we adopt the WeightedMinPair. Proceeding here, we refer to this intermediate algorithm as OnlineRL-N + WeightedMinPair. Despite the superior online fine-tuning performance of this approach, we continue to explore ways to further improve the online learning efficiency by leveraging the ensemble characteristics.
@@ 4.3 Optimistic Exploration
Specifically, we compare three ensemble-based exploration methods. (a) Bootstrapped DQN (Osband et al., 2016) uses ensembles to address some shortcomings of alternative posterior approximation schemes, whose network consists of a shared architecture with N bootstrapped “heads” branching off independently. (b) OAC (Ciosek et al., 2019) proposes an off-policy exploration strategy that adjusts to maximize an upper confidence bound to the critic, obtained from an epistemic uncertainty estimate on the Q-function computed with the bootstrap through Q-ensembles. (c) SUNRISE (Lee et al., 2021) presents ensemble-based weighted Bellman backups that improve the learning process by re-weighting target Q-values based on uncertainty estimates.
The results of different exploration methods is presented in Figure 4. Among them, OnlineRL-N + WeightedMinPair + SUNRISE achieves the highest aggregated return. Consequently, we turn OnlineRL-N + WeightedMinPair + SUNRISE into our final ensemble-based framework E2O. Algorithm 1 summarizes the offline and online procedures of E2O. Note that as many offline RL algorithms can integrate ensemble technique in Q-functions, E2O can thus serve as a common plugin. We will further show the plug-and-play character of E2O by applying OfflineRL-N normal-$\to$ OnlineRL-N + WeightedMinPair + SUNRISE on different offline RL algorithms in the experiments. For a comprehensive view of the detailed results of this section, appending the combination of RandomMinPair and different exploration methods, please refer to Appendix D.4.
@@ 5 Experiments
In this section, we present the empirical evaluations of our E2O framework. We begin with locomotion tasks from D4RL (Fu et al., 2020) to measure the training stability, learning efficiency and final performance of E2O by comparing it with several state-of-the-art offline-to-online RL methods. Additionally, we evaluate E2O on more challenging navigation tasks to verify its versatility.
@@ 5.1 Locomotion Tasks
We first evaluate our E2O framework on MuJoCo (Todorov et al., 2012) locomotion tasks, i.e., HalfCheetah, Walker2d, and Hopper from the D4RL benchmark suite (Fu et al., 2020). To demonstrate the applicability of E2O on various suboptimal datasets, we use three dataset types: medium, medium-replay, and medium-expert. Specifically, medium datasets contain samples collected by a medium-level policy, medium-replay datasets include all samples encountered while training a medium-level agent from scratch, and medium-expert datasets consist of samples collected by both medium-level and expert-level policies. We pre-train the agent for 1M training steps in the offline phase and perform online fine-tuning for 250K environmental steps. Additional experimental details can be found in Appendix C.
Comparative Evaluation
@@
    AWAC (Nair et al., 2020): an offline-to-online RL method that forces the policy to imitate actions with high advantage estimates in the dataset.
  • $\bullet$
@@
    Balanced Replay (Lee et al., 2022): an offline-to-online RL method that trains an additional network to prioritize samples in order to effectively use new data as well as near-on-policy samples in the offline dataset.
  • $\bullet$
@@
    PEX (Zhang et al., 2023): a recent offline-to-online RL method utilizing an offline policy within a policy set, expanding it with additional policies, and constructing a categorical distribution based on their values at the current state to select the final action.
  • $\bullet$
@@
    IQL (Kostrikov et al., 2021): a representative RL algorithm demonstrating superior offline performance and enabling seamless online fine-tuning through direct parameter transfer.
  • $\bullet$
@@
    SAC (Haarnoja et al., 2018): a SAC agent trained from scratch. This baseline highlights the benefit of offline-to-online RL, as opposed to fully online RL, in terms of learning efficiency.
  • $\bullet$
@@
    Scratch: training SAC-N + WeightedMinPair + SUNRISE online from scratch without offline pre-training, as opposed to our E2O framework.
@@
Figure 5 shows the performance of the E2O-CQL method (E2O instantiated on CQL) and baseline methods during the online fine-tuning phase. Compared with pure online RL methods such as SAC and Scratch, E2O-CQL starts with a well-performed policy and learns quickly and stably, proving the benefits of offline pre-training. For offline RL methods, IQL shows limited improvement as complete pessimistic training is no longer suitable for online fine-tuning, while E2O-CQL displays fast fine-tuning. Among other offline-to-online RL methods, the performance of AWAC is limited by the quality of the dataset due to the operation of training its policy to imitate actions with high advantage estimates, resulting in slow improvement during the online phase. While Balanced Replay can attain performance second only to E2O-CQL on some datasets, it also suffers from unstable training. PEX experiences severe performance drop during the initial stage of online fine-tuning on several datasets as the new policies trained in the online phase are random at the beginning and have a negative impact on the training stability. Overall, our E2O-CQL method outperforms the baseline methods in terms of training stability, learning efficiency and final performance on most tasks.
@@ 5.2 Navigation Tasks
We further verify the effectiveness of E2O on D4RL navigation task Antmaze (Fu et al., 2020) by integrating another offline RL algorithm LAPO (Chen et al., 2022). In detail, we specialize E2O as LAPO-N + WeightedMinPair + SUNRISE, i.e., E2O-LAPO. For the Antmaze task, we consider three types of mazes: umaze, medium and large mazes, and two data compositions: play and diverse. The data compositions vary in their action coverage of different regions of the state space and the sub-optimality of the behavior policy.
Comparative Evaluation
Since Antmaze is a more challenging task, most offline RL methods struggle to achieve satisfactory results in the offline phase, we only compare our E2O-LAPO method on this task with two effective baseline methods, IQL and PEX.
Figure 6 presents the performance of E2O-LAPO and baseline methods during the online fine-tuning phase. First, LAPO demonstrates better offline performance than IQL, providing a higher starting point for the online phase, especially in the umaze and medium maze environments where it almost reaches the performance ceiling. In the online stage, IQL shows slower asymptotic performance due to offline policy constraints. Building upon IQL, PEX enhances the degree of exploration by incorporating additional new policies trained from scratch, but the strong randomness of these policies in the early online stage causes performance drop. Note that although both IQL and PEX share the same starting point, PEX exhibits more severe performance drop on most tasks. By employing our proposed E2O framework, we demonstrate that E2O-LAPO can not only enhance the offline performance, but also facilitate stable and rapid performance improvement while maintaining the offline performance without degradation. This approach enables the offline agent to quickly adapt to the real-world environment, providing efficient and effective online fine-tuning. Additionally, we directly leverage LAPO with two Q networks for offline-to-online training and use the comparison with our E2O-LAPO method to further verify the effectiveness of our E2O framework. The results including some ablation studies can be found in Appendix D.
@@ 6 Conclusions and Limitations
In this work, we have demonstrated that Q-ensembles can be efficiently leveraged to alleviate unstable training and performance drop, and serve as a more flexible constraint method for online fine-tuning in various settings. Based on this observation, we propose Ensemble-based Offline-to-Online (E2O) RL Framework, which enables many pessimistic offline RL algorithms to perform optimistic online fine-tuning and improve their performance efficiently while maintaining stable training process. The proposed framework is straightforward and can be combined with many existing offline RL algorithms. We instantiate E2O with different combinations and conducted experiments on a wide range of tasks to demonstrate its effectiveness.
@@
Despite the promising results, there are some limitations to our work that should be acknowledged. First, although E2O is designed to be a flexible plugin for various offline RL algorithms, it may require further modifications or fine-tuning to achieve optimal performance in different contexts. For instance, adjusting the weight coefficient of the BC item may result in better online fine-tuning performance for TD3+BC (Fujimoto & Gu, 2021). Second, the computational cost of using ensembles and uncertainty estimates may limit the scalability of E2O to large-scale problems. Future work could investigate ways to reduce the computational overhead by using deep ensembles (Fort et al., 2019) or ensemble distillation (Hinton et al., 2015) while maintaining the performance benefits of ensembles and uncertainty estimation by using Bayesian compression (Louizos et al., 2017) or variational approximations (Kingma & Welling, 2013). These methods could make E2O more scalable and practical for large-scale problems and real-world applications, enabling the development of more efficient and reliable offline-to-online RL systems.
Appendix A Pseudocode
Algorithm 1 E2O: Ensemble-based Offline-to-Online RL Framework

### v4

@@ 3.2 Loosing Pessimism
Specifically, we compare several Q-target computation methods. (a) MinQ is what we use in OnlineRL-N, where the minimum value of all the Q-networks in the ensemble is taken to compute the target. (b) MeanQ leverages the average of all the Q-values to compute the target. (c) REM is a method originally proposed to boost performance of DQN in the discrete-action setting, which uses the random convex combination of Q-values to compute the target Agarwal et al. (2020). It is similar to ensemble average (MeanQ), but with more randomization. (d) RandomMinPair uses a minimization over a random subset 2 of the $N$ Q-functions, which is proposed in prior methods Chen et al. (2021). (e) WeightedMinPair computes the target as the expectation of all the RandomMinPair targets, where the expectation is taken over all $N$-choose-2 pairs of Q-functions. RandomMinPair can be considered as a uniform-sampled version of WeightedMinPair.
Fig. 3 presents the results of using different Q-target computation methods in the online phase based on OnlineRL-N. With MinQ, which is originally used in OnlineRL-N, as the bound, both MeanQ and REM exhibit poor performance, while RandomMinPair and WeightedMinPair outperform the other candidates with their efficient and stable online learning process. As the WeightedMinPair method is more stable on many datasets than the RandomMinPair method, we adopt the WeightedMinPair. Proceeding here, we refer to this intermediate algorithm as OnlineRL-N + WeightedMinPair. Despite the superior online fine-tuning performance of this approach, we continue to explore ways to further improve the online learning efficiency by leveraging the ensemble characteristics.
@@ 3.3 Optimistic Exploration
Specifically, we compare three ensemble-based exploration methods. (a) Bootstrapped DQN Osband et al. (2016) uses ensembles to address some shortcomings of alternative posterior approximation schemes, whose network consists of a shared architecture with N bootstrapped “heads” branching off independently. (b) OAC Ciosek et al. (2019) proposes an off-policy exploration strategy that adjusts to maximize an upper confidence bound to the critic, obtained from an epistemic uncertainty estimate on the Q-function computed with the bootstrap through Q-ensembles. (c) SUNRISE Lee et al. (2021) presents ensemble-based weighted Bellman backups that improve the learning process by re-weighting target Q-values based on uncertainty estimates.
The results of different exploration methods is presented in Fig. 4. Among them, OnlineRL-N + WeightedMinPair + SUNRISE achieves the highest aggregated return. Consequently, we turn OnlineRL-N + WeightedMinPair + SUNRISE into our final ensemble-based framework ENOTO. Algorithm 1 summarizes the offline and online procedures of ENOTO. Note that as many offline RL algorithms can integrate ensemble technique in Q-functions, ENOTO can thus serve as a common plugin. We will further show the plug-and-play character of ENOTO by applying OfflineRL-N $\to$ OnlineRL-N + WeightedMinPair + SUNRISE on different offline RL algorithms in the experiments. For a comprehensive view of the detailed results of this section, appending the combination of RandomMinPair and different exploration methods, please refer to appendix.
Algorithm 1 ENOTO: ENsemble-based Offline-To-Online RL Framework
@@ 4 Experiments
In this section, we present the empirical evaluations of our ENOTO framework. We begin with locomotion tasks from D4RL Fu et al. (2020) to measure the training stability, learning efficiency and final performance of ENOTO by comparing it with several state-of-the-art offline-to-online RL methods. Additionally, we evaluate ENOTO on more challenging navigation tasks to verify its versatility.
@@ 4.1 Locomotion Tasks
We first evaluate our ENOTO framework on MuJoCo Todorov et al. (2012) locomotion tasks, i.e., HalfCheetah, Walker2d, and Hopper from the D4RL benchmark suite Fu et al. (2020). To demonstrate the applicability of ENOTO on various suboptimal datasets, we use three dataset types: medium, medium-replay, and medium-expert. Specifically, medium datasets contain samples collected by a medium-level policy, medium-replay datasets include all samples encountered while training a medium-level agent from scratch, and medium-expert datasets consist of samples collected by both medium-level and expert-level policies. We pre-train the agent for 1M training steps in the offline phase and perform online fine-tuning for 250K environmental steps. Additional experimental details can be found in the appendix.
Comparative Evaluation.
@@
  • AWAC Nair et al. (2020): an offline-to-online RL method that forces the policy to imitate actions with high advantage estimates in the dataset.
@@
  • BR Lee et al. (2022): an offline-to-online RL method that trains an additional network to prioritize samples in order to effectively use new data as well as near-on-policy samples in the offline dataset.
@@
  • PEX Zhang et al. (2023): a recent offline-to-online RL method utilizing an offline policy within a policy set, expanding it with additional policies, and constructing a categorical distribution based on their values at the current state to select the final action.
  • Cal-QL Nakamoto et al. (2023): a recent offline-to-online RL method learning a conservative value function initialization that underestimates the value of the learned policy from offline data, while also being calibrated, in the sense that the learned Q-values are at a reasonable scale.
@@
  • IQL Kostrikov et al. (2021): a representative RL algorithm demonstrating superior offline performance and enabling seamless online fine-tuning through direct parameter transfer.
@@
  • SAC Haarnoja et al. (2018): a SAC agent trained from scratch. This baseline highlights the benefit of offline-to-online RL, as opposed to fully online RL, in terms of learning efficiency.
@@
  • Scratch: training SAC-N + WeightedMinPair + SUNRISE online from scratch without offline pre-training, as opposed to our ENOTO framework.
@@
Fig. 5 shows the performance of the ENOTO-CQL method (ENOTO instantiated on CQL) and baseline methods during the online fine-tuning phase. Compared with pure online RL methods such as SAC and Scratch, ENOTO-CQL starts with a well-performed policy and learns quickly and stably, proving the benefits of offline pre-training. For offline RL methods, IQL shows limited improvement as complete pessimistic training is no longer suitable for online fine-tuning, while ENOTO-CQL displays fast fine-tuning. Among other offline-to-online RL methods, the performance of AWAC is limited by the quality of the dataset due to the operation of training its policy to imitate actions with high advantage estimates, resulting in slow improvement during the online phase. While BR can attain performance second only to ENOTO-CQL on some datasets, it also suffers from unstable training. PEX exhibits a notable decline in performance during the initial stages of online fine-tuning across various datasets, attributed to the randomness of newly trained policies in the early phase, which negatively affects training stability. Although the original PEX paper does not explicitly address this phenomenon, a meticulous examination of its experimental section reveals that performance drop indeed affects PEX. We contend that the phenomenon of performance drop is a pivotal concern in the domain of offline-to-online RL, warranting significant attention. Turning to the Cal-QL algorithm, while its efficacy is prominently showcased in intricate tasks such as Antmaze, Adroit, and Kitchen, as emphasized in the paper, we note a more subdued performance in traditional MuJoCo tasks. The enhancements during the online phase appear relatively constrained. However, its most salient attribute lies in its exceptional stability, effectively circumventing the issue of performance drop. It is worth noting that the Hopper-medium-expert-v2 dataset represents a special case where most considered offline-to-online RL methods exhibit varying degrees of performance drop, except for Cal-QL, which maintains its offline-stage performance while remaining stable.
It is important to underscore that due to the partial incompleteness of code provided by certain baseline algorithms, our experiments partially rely on publicly available and widely accepted code repositories from GitHub Seno and Imai (2022); Tarasov et al. (2022). Consequently, the experimental results may exhibit slight deviations from the reported outcomes in the original papers, which will be comprehensively detailed in the appendix. Nevertheless, through rigorous comparisons encompassing both the baseline papers’ original performance metrics and the results obtained from our code implementation, our ENOTO method consistently surpasses the baseline approaches in terms of training stability, learning efficiency, and final performance across most tasks. Unfortunately, due to constraints within this text, we can only present the results attained from executing the code, as graphical representations from the source papers cannot be seamlessly incorporated.
@@ 4.2 Navigation Tasks
We further verify the effectiveness of ENOTO on D4RL navigation task Antmaze Fu et al. (2020) by integrating another offline RL algorithm LAPO Chen et al. (2022). In detail, we specialize ENOTO as LAPO-N + WeightedMinPair + SUNRISE, i.e., ENOTO-LAPO. For the Antmaze task, we consider three types of mazes: umaze, medium and large mazes, and two data compositions: play and diverse. The data compositions vary in their action coverage of different regions of the state space and the sub-optimality of the behavior policy.
Comparative Evaluation.
Since Antmaze is a more challenging task, most offline RL methods struggle to achieve satisfactory results in the offline phase, we only compare our ENOTO-LAPO method on this task with three effective baseline methods, IQL, PEX and Cal-QL. Specifically, for the D4RL Antmaze tasks, these methods apply a reward modification following previous works. This modification effectively introduces a survival penalty that encourages the agent to complete the maze as quickly as possible. In the online phase, we maintain the same reward modification as the offline phase during training but keep the rewards unchanged during evaluation.
Fig. 6 presents the performance of ENOTO-LAPO and baseline methods during the online fine-tuning phase. First, LAPO demonstrates better offline performance than IQL, providing a higher starting point for the online phase, especially in the umaze and medium maze environments where it almost reaches the performance ceiling. In the online stage, IQL shows slower asymptotic performance due to offline policy constraints. Building upon IQL, PEX enhances the degree of exploration by incorporating additional new policies trained from scratch, but the strong randomness of these policies in the early online stage causes performance drop. Note that although both IQL and PEX share the same starting point, PEX exhibits more severe performance drop on most tasks. Regarding the Cal-QL algorithm, akin to the outcomes portrayed in the original paper, it demonstrates robust performance in the Antmaze environment, outperforming significantly its MuJoCo counterparts. Notably, it exhibits superior stability and learning efficiency compared to the two baseline methods, IQL and PEX. For our proposed ENOTO framework, we demonstrate that ENOTO-LAPO can not only enhance the offline performance, but also facilitate stable and rapid performance improvement while maintaining the offline performance without degradation. This approach enables the offline agent to quickly adapt to the real-world environment, providing efficient and effective online fine-tuning. Additionally, we directly leverage LAPO with two Q networks for offline-to-online training and use the comparison with our ENOTO-LAPO method to further verify the effectiveness of our ENOTO framework. The results including some ablation studies can be found in the appendix.
@@ 5 Conclusions and Limitations
In this work, we have demonstrated that Q-ensembles can be efficiently leveraged to alleviate unstable training and performance drop, and serve as a more flexible constraint method for online fine-tuning in various settings. Based on this observation, we propose Ensemble-based Offline-to-Online (ENOTO) RL Framework, which enables many pessimistic offline RL algorithms to perform optimistic online fine-tuning and improve their performance efficiently while maintaining stable training process. The proposed framework is straightforward and can be combined with many existing offline RL algorithms. We instantiate ENOTO with different combinations and conducted experiments on a wide range of tasks to demonstrate its effectiveness.
@@
Despite the promising results, there are some limitations to our work that should be acknowledged. First, although ENOTO is designed to be a flexible plugin for various offline RL algorithms, it may require further modifications to achieve optimal performance in different contexts. For instance, adjusting the weight coefficient of the BC item may result in better fine-tuning performance for TD3+BC Fujimoto and Gu (2021). Second, the computational cost of ensembles and uncertainty estimates may limit the scalability of ENOTO to large-scale problems. Future work could investigate ways to reduce the computational overhead by using deep ensembles Fort et al. (2019) or ensemble distillation Hinton et al. (2015), while maintaining the performance by using Bayesian compression Louizos et al. (2017) or variational approximations Kingma and Welling (2013). These methods could make ENOTO more scalable and practical for large-scale problems and real-world applications, enabling the development of more efficient and reliable offline-to-online RL systems.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant Nos. 92370132, 62106172), the Science and Technology on Information Systems Engineering Laboratory (Grant Nos. WDZC20235250409, 6142101220304), and the Xiaomi Young Talents Program of Xiaomi Foundation.
Appendix A Related Works
Offline RL
Offline RL algorithms focus on training RL agents with pre-collected datasets. However, these algorithms face the challenge of distribution shift between the behavior policy and the policy being learned, which can cause issues due to out-of-distribution (OOD) actions sampled from the learned policy and passed into the learned critic. To mitigate this problem, prior methods constrain the learned policy to stay close to the behavior policy via explicit policy regularization Fujimoto et al. [2019]; Wu et al. [2019], via implicit policy constraints Kostrikov et al. [2021]; Chen et al. [2022], by leveraging auxiliary behavioral cloning losses Fujimoto and Gu [2021], by penalizing the Q-value of OOD actions to prevent selecting them Kumar et al. [2020]; An et al. [2021]; Bai et al. [2022]; Yang et al. [2022], or through model-based training with conservative penalties Yu et al. [2020]; Kidambi et al. [2020]. Among the above methods, we choose several representative algorithms such as CQL Kumar et al. [2020] and LAPO Chen et al. [2022] to be the base methods in the offline component of our framework due to their wide applicability and superior performance.
Offline-to-Online RL
Offline-to-online RL refers to the process of improving the well-trained offline policy by incorporating online interactions. Directly applying pre-trained offline policy to the online fine-tuning stage may lead to poor performance due to excess conservatism Nair et al. [2020]; Lee et al. [2022]; Zhao et al. [2022]. To adapt offline RL algorithms to the online environment, several modifications are required. AWAC Nair et al. [2020] is the first algorithm proposed to perform well in the offline-to-online RL setting, which forces the policy to imitate actions with high advantage estimates. AW-Opt Lu et al. [2021] improves upon AWAC by incorporating positive sample filtering and hybrid actor-critic exploration during the online stage. BR Lee et al. [2022] trains an additional neural network to prioritize samples in order to effectively use new data as well as near-on-policy samples from the offline dataset. PEX Zhang et al. [2023] proposes a policy expansion approach for offline-to-online RL, which trains more policies from scratch in the online phase and combine them with the pre-trained offline policy to make decisions jointly. Cal-QL Nakamoto et al. [2023] is a recent offline-to-online RL method learning a conservative value function initialization that underestimates the value of the learned policy from offline data, while also being calibrated, in the sense that the learned Q-values are at a reasonable scale. A concurrent work with us is O3F Mark et al. [2022], which also aims to eliminate pessimism in online learning. However, our research takes a more holistic perspective by connecting the pessimistic offline learning and optimistic online phases through ensemble modeling, proposing a more applicable framework that encapsulates the implementation of O3F as a special case.
Q-Ensembles in RL
Q-Ensemble methods have been widely utilized to enhance the performance of RL Osband et al. [2016]; Fujimoto et al. [2018]; Ciosek et al. [2019]; Chen et al. [2021]; Lee et al. [2021]. TD3 Fujimoto et al. [2018] leverages an ensemble of two value functions and uses their minimum for computing the target value during Bellman error minimization. REDQ Chen et al. [2021] minimizes over a random subset of Q-functions in the target to reduce over-estimation bias. For offline RL, a number of works have extended this to propose backing up minimums or lower confidence bound estimates over larger ensembles Fujimoto et al. [2019]; Wu et al. [2019]; Agarwal et al. [2020]; An et al. [2021]; Bai et al. [2022]; Yang et al. [2022]. In particular, EDAC An et al. [2021] achieves impressive performance by simply increasing the number of Q-networks along with the clipped Q-learning and further proposes to reduce the required number of ensemble networks through ensemble gradient diversification for the purpose of reducing computational cost. PBRL Bai et al. [2022] and RORL Yang et al. [2022] both employ an ensemble of bootstrapped Q-functions for uncertainty quantification and perform pessimistic updates to penalize Q functions with high uncertainties. Recent work Ghasemipour et al. [2022] advocates for using independently learned ensembles without sharing target values and optimizing a policy based on the lower confidence bound of predicted action values. In our work, we discover that Q-ensembles are highly effective in addressing the performance degradation that occurs in the offline-to-online setting, and we advocate for using Q-ensembles for both offline and online algorithms to achieve steady and sample-efficient offline-to-online RL.
Exploration Mechanisms
Various exploration approaches have been proposed to accelerate the efficiency of online training in recent years. These methods can typically be divided into two main categories following Hao et al. [2023]: uncertainty-oriented exploration and intrinsic motivation-oriented exploration. The former employs heuristic design to formulate various intrinsic motivations for exploration, based on factors such as visitation count Ostrovski et al. [2017]; Tang et al. [2017], curiosity Savinov et al. [2018], information gain Houthooft et al. [2016], etc. Uncertainty-oriented exploration methods adopt the principle of optimism in the face of uncertainty and use Q-ensembles to encourage agents to explore areas with higher uncertainty Ecoffet et al. [2021]; Lee et al. [2021]; Liu et al. [2024]. In this paper, we investigate uncertainty-based exploration mechanisms under the unified framework of Q-ensembles to enhance the performance during the online fine-tuning phase.